OpenStudy (anonymous):

Am I correct? Find 11π / 18 radians to degrees. My answer: 11π/18 * 180/π = 110π/π =110 degrees Am I correct? Thank you.

OpenStudy (mertsj):

Yes

OpenStudy (anonymous):

Thank you.

OpenStudy (mertsj):

yw

OpenStudy (amistre64):

since we have a pi, we have to divide off the pi; so yes

OpenStudy (anonymous):

\[\frac{11 \pi }{18} \frac{180}{\pi }=110 \]